# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v3

@@ 2.5.2 Generative View Synthesis
Earlier works primarily used transformer-based autoregressive models [124, 149]. For example, GFVS [150] approaches novel view synthesis by sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively with a transformer. ViewFormer [151] further extends single-view NVS of GFVS to multiview NVS through a proposed branching attention. Recently, latent diffusion models [152] have been widely used in novel view synthesis due to their ability in generating high-resolution images. These models encode inputs into a latent space using a pretrained variational autoencoder and perform diffusion within the latent space. We categorize them into latent image diffusion models and video diffusion models.
@@
Image Diffusion Models. Zero-1-to-3 [147] leverages the latent diffusion model [152] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 7(a). ZeroNVS [153] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis, with an additional field of view as camera parameters to solve the scale ambiguity. However, these methods still face challenges in generating consistent novel views. SyncDreamer [154] addresses this by initializing a diffusion model with Zero-1-to-3 weights and modeling the joint distribution of multiview images. Zero123++[12] arranges six views in a single image for improved multiview representation modeling. ViewDiff [155] also exploits diffusion priors, with an emphasis on multiview consistency in real-world data. Consistent123 [156] combines Zero-1-to-3 and stable diffusion to provide case-aware diffusion priors to ensure multiview-consistency, while ConsistNet [157] performs parallel viewpoint-specific diffusion and aligns the generated images accordingly to enforce multiview geometric consistency. MVDream [88] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. With similar multiview diffusion architectures, CAT3D [158] & SEVA [159] and CAT4D [160] substantially advance the performance for static and dynamic view synthesis.
@@
Video Diffusion Models. Video diffusion models [161, 162] have achieved impressive performance in video synthesis and are believed to implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging video diffusion priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [148], shown in Fig. 7(b), uses the generative prior of large pretrained video diffusion models [162] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Instead of directly using point clouds as condition, ViewCrafter [8] renders point cloud into images as conditions of the video diffusion model [162] to enable consistent and accurate novel view synthesis. MultiDiff [163] leverages a single reference image and a predefined target camera trajectory as conditions of diffusion models, with depth cues to encourage consistent novel view synthesis. More recently, DifFusion3D+ [164], SpatialCrafter [165], and GenFusion [166] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
The preceding methods section highlights the core architectures of feed-forward models, which also include the corresponding basic tasks of novel view synthesis or static 3D reconstruction. In contrast, this section extends beyond the basic tasks and explores the wide variety of applications enabled by feed-forward models.  [^1] We do not cover tasks and applications centered on depth, as this is an independent research line, even though depth is sometimes an output of feed-forward models.
@@ 3.1 Camera Pose Estimation
Feed-forward models that predict pointmaps enable efficient recovery of camera parameters without the need for explicit multi-stage SfM pipelines. DUSt3R demonstrates how pose estimation tasks can be naturally derived from pointmap output. Specifically, camera intrinsics such as the focal length can be solved by minimizing reprojection errors of pointmaps. The relative camera motion between two views can be recovered either by first performing pixel matching and intrinsics estimation to compute the essential matrix [167], or more directly by aligning the predicted pointmaps across views through Procrustes alignment [168], which yields rotation and translation up to scale. To improve robustness against noise and outliers, pointmap alignment can also be integrated with PnP-RANSAC [169, 167, 170].
Instead of relying on pointmaps to derive camera parameters, another line of research [171, 3, 2] aims to predict camera parameters directly in a feed-forward manner. Typically, a learnable camera token is prepended to the image tokens, which allows for interaction with other views to capture image-level information, e.g., camera motions. On the other hand, most of them anchor their pose estimation to a reference viewpoint like the first frame, which is an inductive bias that can lead to inferior results when the reference is suboptimal. $\pi^{3}$ [172] instead proposes to estimate the relative camera pose between views, achieving superior performance of camera pose estimation and benefiting other associated tasks such as pointmap estimation.
@@ 3.2 Pose-Free 3D Reconstruction & View Synthesis
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [173] uses a single-view feed-forward NeRF combined with optical flow to estimate poses and fuse multiview point maps for NeRF-based 3D reconstruction. CoPoNeRF [174] builds and refines 4D correlation maps from image pairs to estimate flow and poses, enabling color and depth rendering.
@@
To extend these approaches to 3DGS, GGRt [93] employs PixelSplat [5] for predicting viewpoint-specific 3D Gaussian maps and introduces a pose estimation module that jointly optimizes camera poses alongside Gaussian predictions. PF3plat [175] proposes a coarse-to-fine strategy, estimating depth, confidence, and camera poses from input images to guide the prediction of 3D Gaussians.
@@
Additionally, several methods build upon DUSt3R [1] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [62] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [59] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously, and significantly enhancing reconstruction quality. Conversely, CUT3R [3] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content.
@@
Based on pointmap reconstruction, several methods have further developed high-quality novel view synthesis through 3D Gaussian reconstruction. Splatt3R [76] extends DUSt3R by adding a Gaussian head decoder that predicts Gaussian parameters directly from image pairs. LSM [176] similarly integrates a Gaussian head and further incorporates semantic embeddings from input images to augment anisotropic Gaussian predictions. NoPosplat [102], after integrating a Gaussian head, performs full-parameter training to predict 3D Gaussians in a canonical space without relying on ground-truth camera poses or depth. PREF3R [105], based on Spann3R, also adds a Gaussian head to achieve 3D Gaussian predictions. SmileSplat [103], another Spann3R derivative, opts to predict Gaussian surfels instead of traditional 3D Gaussians. SelfSplat [104] integrates DUSt3R-based Gaussian predictions with self-supervised depth and pose estimation, jointly predicting depth, camera poses, and Gaussian attributes in a unified neural network. Lastly, FLARE [171] incorporates additional modules for pose estimation and global geometry projection, facilitating alignment of DUSt3R-based network token outputs.
@@
Recent research has also explored pose-free feed-forward approaches at the object level. FORGE [177] transforms per-view voxel features into a shared space using estimated relative camera poses and fuses them into a neural volume for rendering. LEAP [178] selects a canonical view from the input images, defines the neural volume in its local camera coordinate system, and reconstructs a radiance field by iteratively updating the volume via multiview encoding and a 2D-to-3D mapping module. PF-LRM [44] jointly predicts a triplane NeRF and relative poses from sparse unposed images, supervising reconstruction with rendering losses and refining poses via a differentiable PnP solver. MVDiffusion++[179] enables 3D consistency across views through 2D self-attention and view dropout, enabling dense and high-resolution synthesis without explicit pose supervision. SpaRP [180] pushes further by integrating sparse, unposed views into a composite image, which is then processed by a finetuned 2D diffusion model to enable both pose estimation and textured mesh reconstruction.
@@ 3.3 Dynamic 3D Reconstruction & View Synthesis
Compared to the basic task of static scene reconstruction, dynamic scene reconstruction [181] poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@
Seminal work on monocular depth estimation methods learned to predict temporally consistent depth video using temporal attention layers [182] and generative priors [183, 184]. Though they demonstrate pleasure 3D points on camera space, they fail to provide global scene geometry due to the lack of camera pose estimation.
@@
To jointly resolve pose and obtain a point cloud in canonical space, Robust-CVD [185] and CasualSAM [186] integrate a depth prior with geometric optimization to estimate a smooth camera trajectory, as well as a detailed and stable depth and motion map reconstruction. Most recently, MegaSaM [187] further improves pose and depth accuracy by combining the strengths of several prior works, including DROID-SLAM [188], optical flow [189], and a monocular depth estimation model [190], leading to results with previously unachievable quality.
@@
Alternatively, instead of taking advantage of monocular prior models, some methods aim to train a dynamic 3D model from multiview 3D reconstruction models, e.g., DUSt3R [1]. MonST3R [191] estimates the pointmap at each timestep and processes them using a temporal sliding window to compute pairwise pointmap for each frame pair with MonST3R and optical flow from an off-the-shelf method. These intermediates then serve as inputs to optimize a global point cloud and per-frame camera poses and intrinsics. Video depth can be directly derived from this unified representation. To speed up the optimization process in MonST3R, DAS3R [192] trains a dense prediction transformer [193] for motion segmentation inference and models the static scene as Gaussian splats with dynamics-aware optimization, allowing for more accurate background reconstruction results. Recent work CUT3R [3] fine-tunes MonST3R [191] on both static and dynamic datasets, achieving feedforward reconstruction but without predicting dynamic object segmentation, thereby entangling the static scene with dynamic objects. Although effective, these methods require costly training on diverse motion patterns to generalize well. In contrast, Easi3R [194] takes a different path, exploring a training-free and plug-and-play adaptation that enhances the generalization of DUSt3R variants for dynamic scene reconstruction, achieving accurate dynamic region segmentation, camera pose estimation, and 4D dense point map reconstruction at almost no additional cost on top of DUSt3R. Driv3R [65] further enables dynamic 3D reconstruction in large-scale autonomous driving scenarios by introducing a memory mechanism that supports efficient temporal integration. Besides, it also eliminates the global alignment optimization to reduce computational cost.
@@
In addition to pointmap-based dynamic scene reconstruction, several recent methods based on 3D Gaussian Splatting (3DGS) have also been proposed for feed-forward dynamic reconstruction. L4GM [195] proposes the first 4D reconstruction model that produces animated objects from single-view videos using per-frame 3DGS representation. 4D-LRM [196] builds upon a transformer-based large reconstruction model, leveraging data-driven training for dynamic object reconstruction. It draws inspiration from 4D Gaussian Splatting [197] and reconstructs dynamic objects as anisotropic 4D Gaussian clouds. While prior works focus on dynamic object reconstruction, BulletTimer [6] introduces the first feed-forward model for dynamic scene reconstruction. Building on GS-LRM [81], it incorporates a bullet-time embedding into the input frames and aggregates information across all context frames, enabling feed-forward 3D Gaussian Splatting reconstruction at a specific timestamp. In addition, DGS-LRM [198] presents the first feed-forward deformable 3D Gaussian prediction from monocular videos using a transformer-based LRM, while 4DGT [199] extends transformer-based modeling to 4D Gaussian prediction from real-world monocular videos, demonstrating scalability in real-world settings.
@@
Another line of research focuses on leveraging video pre-trained models for point map prediction by modeling 3D scenes as geometry videos. These approaches utilize diffusion models to learn the joint distribution of multiview RGB and geometric frames. A geometry video consists of standard RGB channels augmented with geometry channels, which encode structural information such as depth [200], XYZ coordinates [201], color point rendering [202, 203], or a combination of point-depth-ray maps [204]. Notably, Aether [205] presents a unified framework that takes as input both image and action latents — such as ray maps — and produces predictions for images, actions, and depth. By flexibly combining different input conditions, Aether successfully achieved 4D dynamic reconstruction from video-only input, image-to-video generation from a single image, and camera-conditioned video synthesis given an image and a camera trajectory.
@@
To enable 3D point tracking, Stereo4D [206] proposes a dynaDUSt3R architecture by incorporating a motion head for scene flow prediction. They use stereo videos from the Internet to create a dataset of more than 100,000 real-world 4D scenes with metric scale and long-term 3D motion trajectories for training. Instead of predicting point map and flow map at reference and target viewpoints, St4RTrack [207] outputs two point maps of different time steps for the reference view given two dynamic frames. The network is trained by reprojected supervision signals, including 2D trajectories and monocular depth, without the need for direct scene flow annotation. Inspired by ZeroCo [208], D2USt3R [209] establishes dense correspondence of two pointmaps using the cross-attention maps of DUSt3R [1].

### v4

@@ 2.5.2 Generative View Synthesis
Earlier works primarily used transformer-based autoregressive models [VDOV17, ERO21]. For example, GFVS [REO21] approaches novel view synthesis by sampling target images from a learned distribution conditioned on a source image and camera transformation, where the distribution is modeled autoregressively with a transformer. ViewFormer [KDSB22] further extends single-view NVS of GFVS to multiview NVS through a proposed branching attention. Recently, latent diffusion models [RBL22] have been widely used in novel view synthesis due to their ability in generating high-resolution images. These models encode inputs into a latent space using a pretrained variational autoencoder and perform diffusion within the latent space. We categorize them into latent image diffusion models and video diffusion models.
@@
Image Diffusion Models. Zero-1-to-3 [LWVH23] leverages the latent diffusion model [RBL22] pretrained for text-to-image generation and replaces text embedding with relative camera poses as conditioning to achieve novel view synthesis as illustrated in Fig. 6(a). ZeroNVS [SLS24] extends Zero-1-to-3 to achieve single-view scene-level novel view synthesis, with an additional field of view as camera parameters to solve the scale ambiguity. However, these methods still face challenges in generating consistent novel views. SyncDreamer [LLZ23] addresses this by initializing a diffusion model with Zero-1-to-3 weights and modeling the joint distribution of multiview images. Zero123++[SCZ23] arranges six views in a single image for improved multiview representation modeling. ViewDiff [HBM24] also exploits diffusion priors, with an emphasis on multiview consistency in real-world data. Consistent123 [WYW23] combines Zero-1-to-3 and stable diffusion to provide case-aware diffusion priors to ensure multiview-consistency, while ConsistNet [YCD24] performs parallel viewpoint-specific diffusion and aligns the generated images accordingly to enforce multiview geometric consistency. MVDream [SWY23] proposes a multiview diffusion model that leverages both 2D and 3D data, combining the generalizability of 2D diffusion models with the consistency of 3D renderings. With similar multiview diffusion architectures, CAT3D [GHH24] & SEVA [ZGV25] and CAT4D [WGP25] substantially advance the performance for static and dynamic view synthesis.
@@
Video Diffusion Models. Video diffusion models [HSG22, XXZ24] have achieved impressive performance in video synthesis and are believed to implicitly capture 3D structures. Building on this capability, recent approaches have explored leveraging video diffusion priors to generate multiview images for high-quality 3D reconstruction. For example, ReconX [LSW24], shown in Fig. 6(b), uses the generative prior of large pretrained video diffusion models [XXZ24] to synthesize novel views. It encodes extracted point clouds as 3D structural conditions, ensuring multiview consistency in the generated novel views. Instead of directly using point clouds as condition, ViewCrafter [YXY24] renders point cloud into images as conditions of the video diffusion model [XXZ24] to enable consistent and accurate novel view synthesis. MultiDiff [MSR24] leverages a single reference image and a predefined target camera trajectory as conditions of diffusion models, with depth cues to encourage consistent novel view synthesis. More recently, DifFusion3D+ [WZT25], SpatialCrafter [ZXG25], and GenFusion [WXH25] have bridged reconstruction and generation by using video diffusion models as scene reconstruction refiners, enabling both artifact removal and scene content expansion.
@@ 3 Tasks & Applications
The preceding methods section highlights the core architectures of feed-forward models, which also include the corresponding basic tasks of novel view synthesis or static 3D reconstruction. In contrast, this section extends beyond the basic tasks and explores the wide variety of applications enabled by feed-forward models. Notably, we do not cover tasks and applications centered on depth, as this is an independent research line, even though depth is sometimes an output of feed-forward models.
@@ 3.1 Camera Pose Estimation
Feed-forward models that predict pointmaps enable efficient recovery of camera parameters without the need for explicit multi-stage SfM pipelines. DUSt3R demonstrates how pose estimation tasks can be naturally derived from pointmap output. Specifically, camera intrinsics such as the focal length can be solved by minimizing reprojection errors of pointmaps. The relative camera motion between two views can be recovered either by first performing pixel matching and intrinsics estimation to compute the essential matrix [HZ03], or more directly by aligning the predicted pointmaps across views through Procrustes alignment [LH99], which yields rotation and translation up to scale. To improve robustness against noise and outliers, pointmap alignment can also be integrated with PnP-RANSAC [FB81, HZ03, LMNF09].
Instead of relying on pointmaps to derive camera parameters, another line of research [ZWX25, WZH25, WCK25] aims to predict camera parameters directly in a feed-forward manner. Typically, a learnable camera token is prepended to the image tokens, which allows for interaction with other views to capture image-level information, e.g., camera motions. On the other hand, most of them anchor their pose estimation to a reference viewpoint like the first frame, which is an inductive bias that can lead to inferior results when the reference is suboptimal. $\pi^{3}$ [WZZ25] instead proposes to estimate the relative camera pose between views, achieving superior performance of camera pose estimation and benefiting other associated tasks such as pointmap estimation.
@@ 3.2 Pose-Free 3D Reconstruction & View Synthesis
The development of feed-forward models has enabled the reconstruction of 3D scenes from unposed images or videos without the need for per-scene optimization. FlowCam [SDTS23] uses a single-view feed-forward NeRF combined with optical flow to estimate poses and fuse multiview point maps for NeRF-based 3D reconstruction. CoPoNeRF [HJS24b] builds and refines 4D correlation maps from image pairs to estimate flow and poses, enabling color and depth rendering.
@@
To extend these approaches to 3DGS, GGRt [LGZ24] employs PixelSplat [CLTS24] for predicting viewpoint-specific 3D Gaussian maps and introduces a pose estimation module that jointly optimizes camera poses alongside Gaussian predictions. PF3plat [HJS24a] proposes a coarse-to-fine strategy, estimating depth, confidence, and camera poses from input images to guide the prediction of 3D Gaussians.
@@
Additionally, several methods build upon DUSt3R [WLC24a] for pose-free 3D reconstruction. DUSt3R itself, as a pioneering feed-forward method, utilizes a transformer-based architecture to regress 3D point maps directly from image pairs. Spann3R [WA24] augments DUSt3R with a spatial memory network, allowing multiview inputs and improving efficiency by eliminating global alignment. However, Spann3R’s sequential processing introduces error accumulation in reconstruction. Fast3R [YSL25] overcomes this limitation by introducing a global fusion transformer, processing multiple views simultaneously, and significantly improving reconstruction quality. In contrast, CUT3R [WZH25] refines sequential reconstruction by maintaining and incrementally updating a persistent internal state that encodes scene content.
@@
Based on pointmap reconstruction, several methods have further developed high-quality novel view synthesis through 3D Gaussian reconstruction. Splatt3R [SZLP24] extends DUSt3R by adding a Gaussian head decoder that predicts Gaussian parameters directly from image pairs. LSM [FZC24a] similarly integrates a Gaussian head and further incorporates semantic embeddings from input images to augment anisotropic Gaussian predictions. NoPosplat [YLX24], after integrating a Gaussian head, performs full-parameter training to predict 3D Gaussians in a canonical space without relying on ground-truth camera poses or depth. PREF3R [CYY24], based on Spann3R, also adds a Gaussian head to achieve 3D Gaussian predictions. SmileSplat [LFTL24], another Spann3R derivative, opts to predict Gaussian surfels instead of traditional 3D Gaussians. SelfSplat [KYP24] integrates DUSt3R-based Gaussian predictions with self-supervised depth and pose estimation, jointly predicting depth, camera poses, and Gaussian attributes in a unified neural network. Lastly, FLARE [ZWX25] incorporates additional modules for pose estimation and global geometry projection, facilitating alignment of DUSt3R-based network token outputs.
@@
Recent research has also explored pose-free feed-forward approaches at the object level. FORGE [JJGZ24] transforms per-view voxel features into a shared space using estimated relative camera poses and fuses them into a neural volume for rendering. LEAP [JJZH23] selects a canonical view from the input images, defines the neural volume in its local camera coordinate system, and reconstructs a radiance field by iteratively updating the volume via multiview encoding and a 2D-to-3D mapping module. PF-LRM [WTB23] jointly predicts a triplane NeRF and relative poses from sparse unposed images, supervising reconstruction with rendering losses and refining poses via a differentiable PnP solver. MVDiffusion++[TCW24] enables 3D consistency across views through 2D self-attention and view dropout, enabling dense and high-resolution synthesis without explicit pose supervision. SpaRP [XLC24] pushes further by integrating sparse, unposed views into a composite image, which is then processed by a finetuned 2D diffusion model to enable both pose estimation and textured mesh reconstruction.
@@ 3.3 Dynamic 3D Reconstruction & View Synthesis
Compared to the basic task of static scene reconstruction, dynamic scene reconstruction [TKB23] poses significant challenges mainly due to the presence of moving objects, changing viewpoints, and temporal variations in scene geometry. Extending feedforward 3D reconstruction for dynamic scenarios mainly involves robust pose estimation to mitigate moving object interference, together with dynamic area segmentation for updating changing environments.
@@
Seminal work on monocular depth estimation methods learned to predict temporally consistent depth video using temporal attention layers [LHS20] and generative priors [HGL25, SYZ25]. Though they demonstrate pleasure 3D points on camera space, they fail to provide global scene geometry due to the lack of camera pose estimation.
@@
To jointly resolve pose and obtain a point cloud in canonical space, Robust-CVD [KRH21] and CasualSAM [ZCL22] integrate a depth prior with geometric optimization to estimate a smooth camera trajectory, as well as a detailed and stable depth and motion map reconstruction. Most recently, MegaSaM [LTC25] further improves pose and depth accuracy by combining the strengths of several prior works, including DROID-SLAM [TD21], optical flow [WLD24], and a monocular depth estimation model [YKH24], leading to results with previously unachievable quality.
@@
Alternatively, instead of taking advantage of monocular prior models, some methods aim to train a dynamic 3D model from multiview 3D reconstruction models, e.g., DUSt3R [WLC24a]. MonST3R [ZHH25] estimates the pointmap at each timestep and processes them using a temporal sliding window to compute pairwise pointmap for each frame pair with MonST3R and optical flow from an off-the-shelf method. These intermediates then serve as inputs to optimize a global point cloud and per-frame camera poses and intrinsics. Video depth can be directly derived from this unified representation. To speed up the optimization process in MonST3R, DAS3R [XTPY24] trains a dense prediction transformer [RBK21] for motion segmentation inference and models the static scene as Gaussian splats with dynamics-aware optimization, allowing for more accurate background reconstruction results. Recent work CUT3R [WZH25] fine-tunes MonST3R [ZHH25] on both static and dynamic datasets, achieving feedforward reconstruction but without predicting dynamic object segmentation, thereby entangling the static scene with dynamic objects. Although effective, these methods require costly training on diverse motion patterns to generalize well. In contrast, Easi3R [CCX25] takes a different path, exploring a training-free and plug-and-play adaptation that enhances the generalization of DUSt3R variants for dynamic scene reconstruction, achieving accurate dynamic region segmentation, camera pose estimation, and 4D dense point map reconstruction at almost no additional cost on top of DUSt3R. To mitigate the conflict between camera pose estimation and geometry reconstruction requires, PAGE-4D [ZWC25] proposes to disentangle static and dynamic information by predicting a dynamic mask, suppressing motion cues for pose estimation while amplifying them for geometry reconstruction. Driv3R [FZD24] further enables dynamic 3D reconstruction in large-scale autonomous driving scenarios by introducing a memory mechanism that supports efficient temporal integration. Besides, it also eliminates the global alignment optimization to reduce computational cost.
@@
In addition to pointmap-based dynamic scene reconstruction, several recent methods based on 3D Gaussian Splatting (3DGS) have also been proposed for feed-forward dynamic reconstruction. L4GM [RXM24] proposes the first 4D reconstruction model that produces animated objects from single-view videos using per-frame 3DGS representation. 4D-LRM [MCY25] builds upon a transformer-based large reconstruction model, leveraging data-driven training for dynamic object reconstruction. It draws inspiration from 4D Gaussian Splatting [YYPZ23] and reconstructs dynamic objects as anisotropic 4D Gaussian clouds. While prior works focus on dynamic object reconstruction, BulletTimer [LRM24] introduces the first feed-forward model for dynamic scene reconstruction. Building on GS-LRM [ZBT24], it incorporates a bullet-time embedding into the input frames and aggregates information across all context frames, enabling feed-forward 3D Gaussian Splatting reconstruction at a specific timestamp. In addition, DGS-LRM [LLW25] presents the first feed-forward deformable 3D Gaussian prediction from monocular videos using a transformer-based LRM, while 4DGT [XLD25] extends transformer-based modeling to 4D Gaussian prediction from real-world monocular videos, demonstrating scalability in real-world settings.
@@
Another line of research focuses on leveraging video pre-trained models for point map prediction by modeling 3D scenes as geometry videos. These approaches utilize diffusion models to learn the joint distribution of multiview RGB and geometric frames. A geometry video consists of standard RGB channels augmented with geometry channels, which encode structural information such as depth [LHL25], XYZ coordinates [MZL25], color point rendering [CZL25, RSH25], or a combination of point-depth-ray maps [JZL25]. Notably, Aether [TZW25] presents a unified framework that takes as input both image and action latents — such as ray maps — and produces predictions for images, actions, and depth. By flexibly combining different input conditions, Aether successfully achieved 4D dynamic reconstruction from video-only input, image-to-video generation from a single image, and camera-conditioned video synthesis given an image and a camera trajectory.
@@
To enable 3D point tracking, Stereo4D [JTL25] proposes a dynaDUSt3R architecture by incorporating a motion head for scene flow prediction. They use stereo videos from the Internet to create a dataset of more than 100,000 real-world 4D scenes with metric scale and long-term 3D motion trajectories for training. Instead of predicting point map and flow map at reference and target viewpoints, St4RTrack [FZW25] outputs two point maps of different time steps for the reference view given two dynamic frames. The network is trained by reprojected supervision signals, including 2D trajectories and monocular depth, without the need for direct scene flow annotation. Inspired by ZeroCo [AKP25], D2USt3R [HAJ25] establishes dense correspondence of two pointmaps using the cross-attention maps of DUSt3R [WLC24a].